A pair of diamond earrings is priced at $325 and is discounted to $150. Find the markdown percentage rate to the nearest whole number.
Answer by stanbon(75887)

A pair of diamond earrings is priced at $325 and is discounted to $150. Find the markdown percentage rate to the nearest whole number.
--------------
Discount quantity = 325-150 = 175
Discount rate = 175/325 =0.5385 = 53.85%
Cheers,
Stan H.